40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 90     Entry time: Fri Aug 19 13:46:55 2016
Author: Alena, Calum, Jon, Liz, Steve, Gabriele 
Type: Facility 
Category: Daily Progress 
Subject: Clean room and vacuum complete 

This morning we installed the clean room curtains and washed them. It turns out that the air filters are supposed to be powered at 277V (?) instead of 115V. So right now the flux is quite low. We are looking into the problem: either replace them with 115V modules or install a small transformer.

We also installed the vacuum chamber on the table and connected all the pumps and gauges. There are no leaks and we could pump down easily the empty chamber. We left for lunch when the pressure was at a few 1e-6 Tor and still going down.

ELOG V3.1.3-